OpenStudy (anonymous):

Simplify the following expression: a + 2c - 5b - c + a - b 2a - 6b + 2c 2a - 6b + c a - 6b + c a - 5b + 2c

OpenStudy (cp9454):

B

iYuko (iyuko):

a + 2c - 5b - c + a - b = a + a - 5b - b + 2c - c = 2a - 6b + c

iYuko (iyuko):

@cp9454 Direct answers are against OS CoC

iYuko (iyuko):

Please refrain from doing it. Thanks!

OpenStudy (cp9454):

i don't care about COC of os, if u have any problem report me

Nnesha (nnesha):

which one is direct answer and which one isn't ?

Nnesha (nnesha):

remember explanation + answer = direct answer explain and let them do some work thanks :-)

iYuko (iyuko):

Got it! ;)

Nnesha (nnesha):

gO_Od!! :-)

OpenStudy (acxbox22):

@cp9454 if u really dont care lets get @Zarkon @ganeshie8 here

Nnesha (nnesha):

?? leave it no need to tag mods

Nnesha (nnesha):

@CatalystDuzIt combine like terms add coefficient of same variables ?

OpenStudy (anonymous):

Hmm?

Nnesha (nnesha):

combine like terms \[\huge\rm \color{green}{a} + 2\color{blue}{c} - 5\color{Red}{b} -\color{blue}{ c} + \color{green}{a} - \color{red}{b}\] add coefficient of a add coefficient of b variables for example a + a can be written as 1a+1a = (1+1)a which is 2a now try b and c terms

OpenStudy (anonymous):

2a -6b+c

Nnesha (nnesha):

so how did you get +c ?

OpenStudy (anonymous):

by combining like terms?

Nnesha (nnesha):

yep but what two terms you combined to get +C ?

OpenStudy (anonymous):

2c and -c

Nnesha (nnesha):

what about -6b? how did you get that ?

OpenStudy (anonymous):

combining -5b and -b

Nnesha (nnesha):

gO_Od to go!

OpenStudy (anonymous):

alright thanks man!